OpenStudy (anonymous):

need help with systems of equations -x+2y=-4 2x-3y=8 :)need to find x and y

OpenStudy (anonymous):

@e.mccormick

OpenStudy (muzzack):

well lets first think about what type of method you are going to be using, there are two types, the substituion and the elimination which one do you want to choose

OpenStudy (e.mccormick):

OK. This time there is no obvious relationship. So, did you want to try substitution again or elimination?

OpenStudy (muzzack):

lets do the elimination because it is lil easier @e.mccormick

OpenStudy (anonymous):

substitution best :)

OpenStudy (e.mccormick):

OK, so do substitution, solve one of the equations for x. The top one is probably easier to do since it already has just 1x. It is a -x, but there is only one of them.

OpenStudy (anonymous):

but I dont know how to solve it for x. I have to multiply it or divide it by a number right?, but im not sure what number

OpenStudy (e.mccormick):

Order of operations in reverse. PEMDAS Paren ( ) Exponents Mult and Div Add and Sub Well, we need to reverse these to un-do things or solve for x: Add and Sub Mult and Div Exponents Paren ( ) \(-x+2y=-4\) So first off, we look for addtion or subtractions to move away from the x. There is the 2y being added, so you must subtract it from both sides.

OpenStudy (anonymous):

so -x=4-2y

OpenStudy (anonymous):

which probably gives me -x=2y

OpenStudy (muzzack):

now divide the -x to make it +x to both sides

OpenStudy (anonymous):

x=2y/-x ?

OpenStudy (e.mccormick):

Can't just drop the 4 like that. It needed to stay. -x=4-2y Then divide out the invisible -1.

OpenStudy (muzzack):

so it would be x=-4+2y

OpenStudy (e.mccormick):

-x means -1x

OpenStudy (somy):

isn't is way easier to do elimination??

OpenStudy (muzzack):

now after that you replace x with the other equation

OpenStudy (e.mccormick):

The number 1 likes to hide. @somy In this case, elimination would be quicker. But someone showed her substitution earlier and messed up, so we are going back over it.

OpenStudy (anonymous):

wait so (x=4+2y)?

OpenStudy (muzzack):

2(-4+2y)-3y=8

OpenStudy (anonymous):

I think I have done substitution for the whole paper so I guess I have to stick to it.

OpenStudy (somy):

oh okay then @e.mccormick :)

OpenStudy (e.mccormick):

Now, where it is at there, you have one variable and numbers. So it is just solve for y.

OpenStudy (muzzack):

yes lets use the order of operations

OpenStudy (anonymous):

so y=16??

OpenStudy (muzzack):

hmm let me double check

OpenStudy (e.mccormick):

You got a sign wrong somewhere.

OpenStudy (muzzack):

\[2(-4+2y)-3y=8\] \[-8+4y-3y=8\] \[-8+y\] \[y=16\]

OpenStudy (e.mccormick):

Ah, Muzzack subbed it in with the wrong sign.

OpenStudy (e.mccormick):

No. That is wrong.

OpenStudy (muzzack):

sorrry

OpenStudy (somy):

oh wait wait @Muzzack y cannot be this

OpenStudy (muzzack):

2(−4-2y)−3y=8 −8-4y−3y=8 −8-7y=8 -7y=16 y=-16/7

OpenStudy (e.mccormick):

\(-x+2y=-4 \) \(-x=-4 - 2y\) \(x=4 + 2y\) So that makes for: \(2x-3y=8\) \(2(4 + 2y)-3y=8\)

OpenStudy (somy):

i think it's better for you @tif49 to watch this and do along http://www.youtube.com/watch?v=it3vYdV_oyc

OpenStudy (e.mccormick):

Muzzack, you are still using the wrong sign....

OpenStudy (anonymous):

i think i trust you guys more than myself watching a math video hahaha

OpenStudy (somy):

lol okay then :D @tif49

OpenStudy (muzzack):

2(4+2y)−3y=8 8+4y-3y=8 8+y=8 y=0

OpenStudy (muzzack):

now replace 0 with the other equation

OpenStudy (muzzack):

2x-3(0)=8

OpenStudy (muzzack):

2x=8 x=4

OpenStudy (muzzack):

(4,0) this is ur answer

OpenStudy (muzzack):

am i right @e.mccormick

OpenStudy (anonymous):

thank you so much guys for your help!!! I really couldnt have done it alone. I have a few more questions, could I tag you in my next posts so you can help me?

OpenStudy (e.mccormick):

And here is what I have had typed up for a bit: \(-x+2y=-4 \) \(2x-3y=8\) Elimination: Since the top has a -1x, this is easy. Multiply by it 2, then add! \(-x+2y=-4 \) \(2(-x+2y)=2(-4) \) \(-2x+4y=-8 \) Now I add that to the other one: \(2x-3y=8\) \(-2x+4y=-8 \) \(2x-2x-3y+4y=8-8 \) \(0x+1y=0 \) The x is now 0x, so eliminated! That is why they call it elimination. And the answer for y happens to be easy now: \(y=0\) If I put that in either of the original equations, I get the answer for x: \(-x+2y=-4 \) \(2x-3y=8\) \(-x+0=-4 \) \(2x-0=8\) \(-x=-4 \) \(2x=8\) \(x=4 \) \(x=4\) OK. Now for substitution: Again, I'll use the easy one to work to start with: \(-x+2y=-4 \) \(-x=-4 - 2y\) \(x=4 + 2y\) So, we know what x is... sort of! (if we knew y was 0 we could see the answer at this point) We take this "solved" x and substitute it into the second equation: \(2x-3y=8\) \(2(4 + 2y)-3y=8\) \(8 + 4y-3y=8\) \(y=8-8\) \(y=0\) From there the rest is the same. All of this checked by a graph: https://www.desmos.com/calculator/xnclxtxagq \(\Huge \ddot \smile\)

OpenStudy (somy):

yup now its same as my answer in elimination lol @Muzzack

OpenStudy (e.mccormick):

That long post I did shows both ways, and you can use that graphing tool linked at the end to check your work. Just type in the calculations for the lines and see what they are.

OpenStudy (somy):

that's super clear now @e.mccormick

OpenStudy (anonymous):

THANKYOU guyss!

OpenStudy (muzzack):

\[\Huge \ddot \smile\]

OpenStudy (muzzack):

LaTeX is awesome